OpenStudy (anonymous):

The first term of an arithmetic sequence is -3 and the fifteenth term is 53. What is the common difference of the sequence?

OpenStudy (kitten_is_back):

4

OpenStudy (anonymous):

ok thnx

OpenStudy (kitten_is_back):

yw

OpenStudy (jango_in_dtown):

Let the common difference be d. Here 1st term=a=-3 now 15th term is given by a+(15-1)d=a+14d=-3+14d by question -3+14d=53 and hence 14d=56 and hence d=4

OpenStudy (leozap1):

Please don't give just the answers, it's against the rules and could get you banned.